Chrisman Corporation makes a product with the following standard costs:

Inputs
Direct materials Direct labor Variable overhead

Standard Quantity or Hours 8.8 pounds
1.4 hours
1.4 hours

Standard Price or Rate $2.00 per pound $20.00 per hour $10.00 per hour

The company reported the following results concerning this product in August.

Actual output
Raw materials used in production
Raw materials purchased
Actual direct labor hours
Actual cost of raw materials purchased Actual cost of direct labor
Actual cost of variable overhead

9,500 units 75,000 pounds 85,000 pounds 4,000 hours $76,000 $70,000 $10,000

The company applies variable overhead on the basis of direct labor hours. The direct materials price variance is computed when the materials are purchased.

a. Compute the materials quantity variance.

b. Compute the materials price variance.

c. Compute the labor efficiency variance.

d. Compute the direct labor rate variance.

e. Compute the variable overhead efficiency variance.

f. Compute the variable overhead rate variance.
Answer #1
a Materials quantity variance 17200 Favorable =2*(75000-9500*8.8)
b Materials price variance 94000 Favorable =76000-(85000*2)
c Labor efficiency variance 186000 Favorable =20*(4000-9500*1.4)
d Direct labor rate variance 10000 Favorable =70000-(4000*20)
e Variable overhead efficiency variance 93000 Favorable =10*(4000-9500*1.4)
f Variable overhead rate variance 30000 Favorable =10000-(4000*10)